In a chemistry class, 12 L of a 12% alcohol solution must be mixed with a 2% solution to get a 14% solution. How many liters of 20% solution are needed?

Answer by checkley71(8403)

I'll assume the 2% should be 20%. The only number that makes any sense.
.12*12+.2x=.14(12+x) or 1.44+.2x=1.68+.14x or .2x-.14x=1.68-1.44
or .06x=.24 or x=.24/.06 or x=4 liters of 20% solution is needed.
PROOF: .12*12+.2*4=.14(12+4) or 1.44+.8=.14*16 or 2.24=2.24.